geerky42 (geerky42):

\(\Large \sqrt{-3}\sqrt{-2} \neq \sqrt{6}\) Why?

OpenStudy (anonymous):

because \[\sqrt{a}\sqrt{b}=\sqrt{ab}\] only if \(a, b>0\)

geerky42 (geerky42):

But why? This is my question.

OpenStudy (anonymous):

if you allow negative numbers under the radical, i.e. if you are living in the world fo complex numbers, then for example \(\sqrt{-9}=3\sqrt{-1}\) or \(3i\)